OpenStudy (anonymous):

Simplify (I got an answer I just need someone to check it) : (-5 + 3i) - (-8 + 2i) I got: 3+i

OpenStudy (nikita2):

(-5 + 3i) - (-8 +2i) = -5 -(-8) + (3-2)i = 3+1i

OpenStudy (anonymous):

does it matter if I put the 1 or not??? Or is it wrong if we don't and thanks a lot!

OpenStudy (nikita2):

it is thee same, but it is better without 1)

OpenStudy (anonymous):

thanks so much!